OpenStudy (anonymous):

how would i create a binomial with a radical in the second term, identify its conjugate and multiply the original binomial and the conjugate?

OpenStudy (anonymous):

\[(a+\sqrt{b})(a-\sqrt{b})=a^2-b\]